• With removal mean u_{z}, we reproduce Hongzhe’s exponential growth; see the red line (single precision)

Figure 1: pvar lnrho profile, showing wiggles on the processor boundaries. These are absent when no mean flow is removed.

Figure 2: pcomp Orange, without any averages removed, and single precision. Red: new run, in single precision, with mean flow removed. Black: Hongzhe’s original run, in double precision, also with mean flow removed. Blue: same as black, but with QKTFM invoked.

Figure 3: pcomp Same, but longer.